On January 1, 2020, Selected Limited purchased 120,000 common shares of Choices Ltd. for $24.50 per share. At the time, Choices had 600,000 shares outstanding. At the time of purchase, Choice's equipment was undervalued by $88,000 and had a remaining useful life of 11 years. Both companies have a December 31 year-end.On May 31, Senator declared and paid a cash dividend of $0.25 per share. On December 31, 2020, Choice reported a net income of $360,000. The shares in Choice were trading at $21.75 per share on that day.

A) Assume that the investment in Choice does not represent significant influence, and that Selected classifies the investment as FV-NI. Prepare the necessary journal entries on the books of Selected to record the above transactions.

B) Assume that the investment in Choice does represent significant influence. Prepare the necessary journal entries on the books of Selected Limited to record the above transactions. (5Marks)
Answer #1

PART A) BY FAIR VALUE METHOD

DATE GENERAL JOURNAL DEBIT CREDIT
, JAN 1 , 2020 Investment in Senator        2,940,000
Cash        2,940,000
(To record purchase of Investment.) (120000*24.50)
MAY 31 Cash              30,000
Dividend Income              30,000
(To record dividend received.) (120000*0.25)
Dec 31 No Journal entry Required for Net income.
Dec 31 Unrealized holding Loss -NI (Net income)            330,000
Fair value Adjustment            330,000
(To record fair value adjustment.) (120000*2.75)
(Decrease in share price = 24.50-21.75 = 2.75

PART B) BY EQUITY METHOD Percentage of ownership acquired (120000/600000) = 20

DATE JOURNAL ENTRY DEBIT CREDIT
Jan 1, 2020 Investment in Senator (associate)       2,940,000
Cash           2,940,000
(To record purchase of significant equity investment.) (120000*24.50)
May 31 Cash             30,000
Investment in Senator (associate)                 30,000
(To record dividend received.) (120000*0.25)
Dec 31 Investment in Senator (associate)             72,000
Income from Equity Investment                 72,000
(To record income from Equity Investment. ) (360000*20%)
Dec 31 Income from equity investment                1,600
Investment in Senator (associate)                   1,600
(To record depreciation on undervalued equipment.) (8000*20%)
(Depreciation = 88000 / 11 = 8000)
Dec 31 No Journal entry Required for Fair Value Adjustment.